OpenStudy (anonymous):

What is the exact value of the expression sqrt(189) - sqrt(21) + sqrt(84) ? Simplify if possible.

OpenStudy (harsimran_hs4):

do you mean 2(sqrt(10k) )* (5k^2)) ?

OpenStudy (anonymous):

\[2\sqrt{10k} \times 5k \sqrt{2} \]

OpenStudy (anonymous):

I have some of it, I think..

OpenStudy (harsimran_hs4):

this can be rewritten as 2 * sqrt(10) * k^(1/2) * 5 * k* sqrt(2) 10*sqrt(10*2) * k^(1/2 + 1) clear till now?

OpenStudy (anonymous):

not clear at all.. my lesson says for me to multiply what's out side the radical first, 2 and 5k to get 10k, then multiply what's in the radical, sqrt{10k} and sqrt{2} to get 20k, then it says to combine, 10k sqrt{20k} and simplify, I just don't know how to simplify it.

OpenStudy (harsimran_hs4):

alright though one and the same thing now for simplification you need to combine the integers and variables(k in this case) so your expression can be written as : 10*k * sqrt(20)* sqrt(k) clear?

OpenStudy (anonymous):

so it would be \[10 \times k \times \sqrt{20} \times \sqrt{k}\] ?

OpenStudy (harsimran_hs4):

yes now 10sqrt(20) * k^(1 + 1/2) for the next step 10sqrt(20)*(k^(3/2)) for the final answer

OpenStudy (anonymous):

can you put it in equation form, cause it's confusing me when your putting it like that.

OpenStudy (harsimran_hs4):

\[10\sqrt{20} k ^{3/2} \]

OpenStudy (harsimran_hs4):

better still \[\sqrt{20} = \sqrt{2 \times 2\times 5} = 2 \sqrt{5}\]

OpenStudy (anonymous):

can you please help me with mine too

OpenStudy (anonymous):

that's not one of the answers.. Do you want me to copy and paste the answers on to here?

OpenStudy (harsimran_hs4):

so final answer \[20\sqrt{5} k ^{3/2}\]

OpenStudy (harsimran_hs4):

@vetdoc does this match?

OpenStudy (anonymous):

I have 20 sqrt(5k) but not the 3/2 part

OpenStudy (harsimran_hs4):

ok i got it how your book is going on with the answer we had reached 10k * sqrt(20k) \[\sqrt{20k} = \sqrt{2 \times 2 \times 5 \times k} = 2 \sqrt{ 5 k}\]

OpenStudy (harsimran_hs4):

so final answer\[10k \sqrt{20k} = 10k \times 2 \sqrt{5k} = 20k \sqrt(5k)\]

OpenStudy (harsimran_hs4):

sqrt(5k) square root of 5k

OpenStudy (anonymous):

okay.. I got it. Thanks

OpenStudy (harsimran_hs4):

great :)